# Omnichain Web The Universal Framework for Streamlined Chain Abstraction and Cross-Layer Interaction

### v1

@@ Abstract
The evolution of the Web3 ecosystem has been hindered by fragmented liquidity and limited interoperability across Layer 1 (L1) and Layer 2 (L2) blockchains, which leads to inefficiencies and elevated costs. Omnichain Web addresses these challenges by introducing a comprehensive framework to unify decentralized networks through its core components: OmniRollups, Proof Network, Ragno Network, and Builder Marketplace. This ecosystem enables seamless cross-chain asset settlement, interoperability, and user-friendly decentralized application (dApp) development, driven by innovative technologies such as modular proof networks and trusted execution environments (TEEs). By integrating advanced zero-knowledge proof systems and compatibility with AI agents, Omnichain Web empowers intent-driven and autonomous functionalities, streamlining liquidity management and user interactions across blockchains. Furthermore, its decentralized marketplace for L1 infrastructure reduces operational overhead and promotes scalable, secure, and efficient cross-chain protocols. As a pioneering solution, Omnichain Web seamlessly connects Web2 and Web3, enabling a holistic and interconnected digital economy.

### v2

@@ Abstract
The Web3 ecosystem is highly fragmented, making it difficult for over a billion Web2 businesses, traditional enterprises, and thousands of AI protocols to integrate seamlessly. With the rapid growth of blockchains, rollups, and app-specific chains, cross-chain interactions remain inefficient, and liquidity is deeply fragmented. AI systems, which thrive on data, lack standardized access to blockchain ecosystems, limiting their ability to function autonomously. Furthermore, intent-based interactions, crucial for AI-driven automation, are constrained by the absence of scalable platforms that enable users and developers to express and execute intents across chains. The current solver ecosystem is centralized, with most of the volume controlled by project teams, due to the lack of developer-friendly tools that address liquidity fragmentation and rebalancing challenges.
To solve this multilayered, multidimensional challenge, Dojima’s Omnichain Web introduces a universal framework that abstracts blockchain complexity and unifies Web2, Web3 and AI systems. OmniRollups, Proof Network, Ragno Network, and Builder Marketplace form its core pillars. The Omni Sequencer ensures atomic, secure execution across chains, while Linera microchains enable AI-driven transaction automation. Ragno Network serves as a decentralized marketplace for L1 infrastructure, while the Proof Network provides cryptographic security for omnichain transactions. The Builder Marketplace introduces a solver-driven execution layer, empowering developers to build and monetize intent-based applications without liquidity constraints. By enabling a composable marketplace at the crossroads of Web2 and Web3, Omnichain Web fosters a seamless flow of data, value, and computation across previously isolated ecosystems.
Omnichain Web is a universal framework that mirrors the internet that unifies Web3 and Web2, enabling a seamless, composable application marketplace. By combining Web2 scale with Web3 decentralization, we are building the foundation for the next wave of adoption.